Question
The equation y= -1777x + 27,153 can be used to predict the number y of gun deaths, in the U.S. x years after 2000, that is, x=0 corresponds to 2003, x=y corresponds to 2004, and so on... Predict the number of gun deaths in 2004 and 2007.
In what year will the number of gun deaths be 11,160?
What number of gun deaths predicted for 2004??? and What number gun deaths predicted for 2007?

Answer by stanbon(48535)

The equation f(x)= -1777x + 27,153 can be used to predict the number y of gun deaths, in the U.S. x years after 2000, that is, x=3 corresponds to 2003, x=4 corresponds to 2004, and so on...
-------------------------------------
Predict the number of gun deaths in 2004 and 2007.
f(4) = -1777*4 + 27153 = 20045
f(7) = -1777*7 + 27153 = 14714
----------------------
In what year will the number of gun deaths be 11,160?
11,160 = -1777x + 27153 ; x=9
Year = 2009
====================
Cheers,
Stan H.